OpenStudy (anonymous):

some one please help! What is the solution to the equation? 1/3x + 1/3 (2x - 15) = 3 1/2 x=

OpenStudy (anonymous):

First you have to simplify the equation so it will be x−5=7/2 After that you add 5 to both sides x=17/2 So x= 8 1/2

OpenStudy (anonymous):

Thank You so much!!